OpenStudy (anonymous):

HELP!!!!!!!! Which ratio is equivalent to 4 : 13? A. 8 : 17 B. 16 : 52 C. 12 : 33 D. 20 : 55

jimthompson5910 (jim_thompson5910):

4:13 is the same as 4/13 let's test choice A 8:17 is the same as 8/17 ------------------------------------------------------- so IF 4:13 is equivalent to 8:17, then this equation must be true 4/13 = 8/17

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

you can cross multiply to check 4/13 = 8/17 4*17 = 13*8 68 = 104 The fact that 68 = 104 is false means that 4/13 = 8/17 is false

jimthompson5910 (jim_thompson5910):

so 4:13 is NOT equivalent to 8:17

jimthompson5910 (jim_thompson5910):

you'll check the other choices that way

OpenStudy (anonymous):

ok ^.^ thx dude or girl wht ever but thx

OpenStudy (anonymous):

i got 2 do about 4 more

OpenStudy (anonymous):

im here to help

OpenStudy (anonymous):

ok =)

OpenStudy (anonymous):

ask away

OpenStudy (anonymous):

Which ratio is not equivalent to 45/120

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i belive it's a or d idk

OpenStudy (anonymous):

its A)

OpenStudy (anonymous):

k =) my first guess

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok it seem tht all my first guesses r right =/